# TaSL: Task Skill Localization and Consolidation for Language Model Continual Learning

### v2

@@ Abstract
Language model continual learning (CL) has recently attracted significant interest for its ability to adapt large language models (LLMs) to dynamic real-world scenarios without retraining. A major challenge in this domain is catastrophic forgetting, where models lose previously acquired knowledge upon learning new tasks. Existing approaches commonly utilize multiple parameter-efficient fine-tuning (PEFT) blocks to acquire task-specific knowledge, yet these methods are inefficient and fail to leverage potential knowledge transfer across tasks. In this paper, we introduce a novel CL framework for language models, named Task Skill Localization and Consolidation (TaSL), which boosts knowledge transfer without depending on memory replay. TaSL initially segregates the model into ‘skill units’ based on parameter dependencies, allowing for more precise control. Subsequently, it employs a novel group-wise skill localization technique to ascertain the importance distribution of skill units for a new task. By comparing this importance distribution with those from previous tasks, we implement a fine-grained skill consolidation strategy that retains task-specific knowledge, thereby preventing forgetting, and updates task-shared knowledge, which facilitates bi-directional knowledge transfer. As a result, TaSL achieves an optimal balance between retaining prior knowledge and excelling in new tasks. TaSL also demonstrates strong generalizability, making it suitable for various base models and adaptable to PEFT methods like LoRA. Furthermore, it offers notable extensibility, supporting enhancements through integration with memory replay techniques. Comprehensive experiments conducted on two CL benchmarks, involving models ranging from 220M to 7B parameters, affirm the effectiveness of TaSL and its variants across different settings.
@@ I Introduction
Figure 1: Conceptual illustration of TaSL. By identifying task-relevant areas across both previously accumulated and current tasks, we can consolidate the task-shared and task-specific parameters to facilitate efficient knowledge transfer (KT) and mitigate catastrophic forgetting (CF).
@@
To address these limitations, we introduce Task Skill Localization and Consolidation (TaSL) [19], a novel CL framework designed to improve KT between tasks without relying on memory replay. Our approach is motivated by recent findings that model parameters contribute unevenly to performance [20]. For instance, the authors in [21] uncovered a core region in LLMs that is crucial for all languages, suggesting that preserving these vital regions can mitigate forgetting. Additionally, findings from [14] indicate that fine-tuning with LoRA often preserves many unnecessary parameter changes, pointing to substantial redundancies within PEFT blocks.
@@
Based on these insights, TaSL facilitates KT by pinpointing and consolidating the importance distribution of model parameters across tasks. Initially, TaSL utilizes a group-wise importance-aware skill localization technique that employs gradient trajectories to identify tiny regions within the parameter space that store crucial knowledge for the current task. By comparing the importance distribution with those of previous tasks, we can then differentiate between task-specific and task-shared regions, as illustrated in Figure 1. The subsequent skill consolidation phase then categorically integrates weights from previous tasks with the current one, efficiently mitigating CF.
@@
In detail, TaSL first reconstructs the model or PEFT block into fine-grained “skill units”. A skill unit refers to a distinct subset of model parameters that encapsulates specific functional capabilities or knowledge relevant to a particular task, such as the Query matrix within the self-attention layer. By operating at this finer granularity, we can localize and consolidate task-specific and shared knowledge within a single PEFT block, rather than adapting a separate PEFT block for each task as in previous works (addressing Limitation 1).
@@
The importance-aware skill localization method employs a new group-wise metric to compute importance scores, effectively quantifying the significance of each skill unit for the current task. Our skill consolidation phase, then based on a fine-grained model averaging strategy, effectively manages different types of knowledge. This stage facilitates forward KT by using previously fine-tuned weights as the starting point for new tasks. For backward KT, we merge knowledge from both current and past tasks into localized task-shared skill units, boosting their effectiveness. To prevent CF, we ensure the integrity of skill units containing prior task-specific knowledge is maintained, safeguarding them from being altered by the learning of new tasks (addressing Limitation 2).
@@
Given the widespread adoption and success of LoRA in fine-tuning LLMs, there is a compelling opportunity to optimize the TaSL framework specifically for LoRA. While the initial TaSL framework has shown promising results, its validation has been limited to a specific sub-task, and its broader applicability on general CL benchmarks remains explored. Additionally, the reliance on first-order gradients for skill localization in TaSL may not yield precise importance localization. The skill consolidation phase also involves many hyperparameters, which could complicate the averaging process.
@@
To address these issues, we propose TasLoRA, a variant of TaSL optimized for LoRA. TasLoRA redesigns the LoRA adapter into new skill units based on parameter dependencies, streamlining knowledge management throughout sequential task learning. It incorporates an orthogonal loss during fine-tuning to ensure distinct latent semantic features are captured within the same LoRA adapter. For skill localization, TasLoRA utilizes a new second-order gradient approximate group-wise metric, enhancing the precision of importance scores. In skill consolidation, it shifts from hard-mask to soft-mask averaging, employing an adaptive technique that flexibly integrates task-specific and shared parameters according to the importance of the skill unit. Additionally, we combine TaSL with memory replay to broaden its application scope, introducing the TaSL-M model. Through extensive experiments on two CL benchmarks with four parameter-level backbones, our TaSL framework and its variants excel in mitigating CF and showcase remarkable capabilities for KT, outperforming SOTA methods.
@@
  • We introduce the Task Skill Localization and Consolidation (TaSL) framework for language model continual learning. By pinpointing and amalgamating task-specific and task-shared knowledge at a granular skill unit level, TaSL ensures effective knowledge transfer and significantly reduces catastrophic forgetting, addressing the shortcomings of previous methodologies.
  • We devise various group-wise skill localization and fine-grained skill consolidation techniques. Notably, our skill localization utilizes either first-order gradients or an innovative second-order gradient approximation for assessing parameter importance. For skill consolidation, we implement strategies ranging from categorical hard-masking to adaptive soft-masking, enhancing the flexibility and precision of knowledge integration.
  • The TaSL framework is distinguished by its robust generalizability and extensibility. Its skill units are designed flexibly, allowing for seamless tailoring to PEFT methods, such as LoRA, optimizing performance across different model architectures. Moreover, integrating TaSL with memory replay enables further performance enhancements and broadens its applicability to diverse scenarios.
  • Extensive evaluation on two CL benchmarks demonstrates the superiority of our TaSL framework and its variants in facilitating knowledge transfer and mitigating catastrophic forgetting, especially in memory-free scenarios. Furthermore, TaSL consistently excels across diverse model sizes (from 220M to 7B), various model architectures (T5 and LLaMA-2), and unseen tasks.
@@ II-A Language Model Continual Learning
Conventional continual learning methods can be categorized into three main types: (i) Regularization-based methods impose explicit constraints to preserve the knowledge of previous tasks by penalizing the variation of each parameter based on its contribution to past tasks [12, 22], such as in the EWC [23] method. However, these methods typically restrict the learning of task-shared knowledge by only considering parameter importance from the perspective of past tasks. This limitation often leads to a suboptimal balance between retaining previous knowledge and excelling at new tasks, as it does not account for the relevance of parameters to both past and current tasks. (ii) Rehearsal-based methods mitigate catastrophic forgetting by either storing old training samples [24, 25] or training generative models to provide pseudo-samples of previous tasks [26, 27], enabling data replay during the learning of new tasks. (iii) Architecture-based methods aim to prevent task interference by dynamically expanding model capacity or assigning dedicated parameter subspaces to each task [28, 29]. While effective at minimizing forgetting, these methods often hinder knowledge transfer (KT) between tasks. Our TaSL framework stands out by bi-directionally analyzing parameter importance across both historical and current tasks, enabling effective KT while mitigating forgetting. To further enhance performance, we extend TaSL with memory replay techniques, resulting in TaSL-M, improving its flexibility and applicability across various scenarios.
Language model continual learning with PEFT methods in the era of LLMs often adopt a parameter isolation strategy. This approach involves using a pipeline to learn and select PEFT blocks for each task [11, 30]. However, as the number of tasks increases, the accumulation of PEFT blocks can become inefficient, and achieving KT between blocks remains challenging. More recently, authors in [31] have explored using the Mixture of Experts (MoE) approach to connect multiple PEFT blocks, but this essentially relies on coarse-grained model averaging. This can lead to overemphasizing unimportant weights, contaminating previously acquired task-specific knowledge, and causing forgetting. To address these challenges in the context of PEFT-based CL, we extend the TaSL framework and introduce TasLoRA. By dividing fine-grained skill units based on parameter dependencies within LoRA, TasLoRA enhances KT efficiency, enables more precise control, and effectively mitigates forgetting.
II-B Skill Localization
@@
Motivated by these findings, we address the challenge of catastrophic forgetting in CL. While [20] introduced the concept of “skill localization” to identify critical parameters in pre-trained language models, their approach requires additional steps to identify and retrain these parameters post-fine-tuning, which impacts efficiency. In contrast, our importance-aware skill localization method leverages trajectory gradients during each task’s training phase to identify parameter importance, significantly reducing processing time. Combined with our skill consolidation strategy, which explicitly targets task-specific and shared parameters, our approach effectively mitigates forgetting and achieves knowledge transfer in CL.
@@ III-A Continual Learning Setup
And we denote $f_{k}$ as the model trained up to task $\mathcal{T}_{k}$, while $\hat{f}_{k}$ represents the model after integrating knowledge from $\hat{f}_{k-1}$ and $f_{k}$ through averaging. Our TaSL framework is designed to tackle a more challenging scenario where the model cannot access any historical data during training [18]. To broaden its applicability, we have augmented TaSL with memory replay. In this enhanced setup, known as TaSL-M, we store a random subset of $\left|\mathcal{M}\right|$ samples from each prior task’s training set $\mathcal{D}_{i}$ in memory $\mathcal{M}_{i}$. The model is then trained jointly on the new task data $\mathcal{D}_{k}$ and the accumulated memory $\mathcal{M}_{<k}$, leveraging this memory replay to improve performance.
@@ IV-A Overview
TaSL includes two key components: (i) Skill Localization, utilizing a group-wise importance metric to precisely identify the importance distribution of parameters across tasks, and (ii) Skill Consolidation, which employs an innovative fine-grained model averaging strategy to integrate model weights from both current and previous tasks for effective knowledge transfer.
Figure 2 presents a detailed overview of our proposed TaSL framework, with the following subsections elaborating on each component and the corresponding extensions.
Figure 2: Left: Depiction of reconstructing LoRA as fine-grained skill units. Right: Overview of TaSL. Step 1: We compute the importance scores of skill units for the current task $k$ using our importance-aware skill localization method during fine-tuning. Step 2: Based on a categorical model averaging strategy, the skill consolidation mechanism merges the model $\hat{f}_{k-1}$, which accumulates knowledge of all previous tasks, with the current task’s model $f_{k}$. This integration is strategically guided by the importance distributions of skill units across different tasks. This iterative process continues with the addition of each new task.
@@ IV-B1 Matrix-Level Skill Unit
In this division strategy, we define skill units as individual matrices in the model, such as the query or key matrices in the self-attention layer or the A matrix in LoRA. This approach aligns with the original TaSL framework and offers the advantage of being model-agnostic. It is suitable for both traditional full-parameter fine-tuning and modern parameter-efficient fine-tuning methods like LoRA, ensuring broad generalizability.
@@ IV-B2 LoRA-Tailored Skill Unit
IV-C Importance-aware Skill Localization
@@ IV-C1 First-Order Gradient-Based Metric
In the original TaSL framework, we defined the importance function $s(\cdot)$ as the magnitude of the gradient-weight product:
@@ IV-C2 Second-Order Gradient Approximation Metric
Algorithm 1 Importance-aware Skill Localization
@@
where $\beta\in[0,1]$ and Norm($\cdot$) normalizes the importance scores to the [0, 1] range, ensuring consistency across models. The initial scores, $\mathcal{I}(\hat{\mathcal{U}}_{1})$, are set equal to $\mathcal{I}(\mathcal{U}_{1})$. Following this, the importance distribution for skill units up to task $\mathcal{T}_{k-1}$ is combined with that of the current task, $\mathcal{T}_{k}$, to facilitate the skill consolidation process.
IV-D Skill Consolidation
Following skill localization, the next critical phase is skill consolidation, where knowledge within each skill unit is integrated into a cohesive framework. This process utilizes a sophisticated model averaging approach that accounts for various factors to optimize task performance.
@@ IV-D Skill Consolidation
IV-D1 Static Weighted Consolidation
In the original TaSL framework, this consolidation strategy initiates by establishing importance thresholds $\delta$ through quantiles, selecting the top 20% of skill units based on importance scores. A skill unit $u_{i}^{k}$ is considered important (denoted as $(u_{i}^{k})^{+}$) if its score $\mathcal{I}(u_{i}^{k})$ exceeds $\delta_{k}$, and unimportant ($(u_{i}^{k})^{-}$) otherwise.
The static weighted consolidation strategy tailors parameter integration for each skill unit, according to its importance across different tasks. The method is defined as follows:

### v3

# KlF: Knowledge Localization and Fusion for Language Model Continual Learning

@@ Abstract
Language model continual learning (CL) has recently attracted significant interest for its ability to adapt large language models (LLMs) to dynamic real-world scenarios without retraining. A major challenge in this domain is catastrophic forgetting, where models lose previously acquired knowledge upon learning new tasks. Existing approaches commonly utilize multiple parameter-efficient fine-tuning (PEFT) blocks to acquire task-specific knowledge, yet these methods are inefficient and fail to leverage potential knowledge transfer across tasks. In this paper, we introduce a novel CL framework for language models, named Knowledge Localization and Fusion (KlF), which boosts knowledge transfer without depending on memory replay. KlF initially segregates the model into ‘skill units’ based on parameter dependencies, allowing for more precise control. Subsequently, it employs a novel group-wise knowledge localization technique to ascertain the importance distribution of skill units for a new task. By comparing this importance distribution with those from previous tasks, we implement a fine-grained knowledge fusion strategy that retains task-specific knowledge, thereby preventing forgetting, and updates task-shared knowledge, which facilitates bi-directional knowledge transfer. As a result, KlF achieves an optimal balance between retaining prior knowledge and excelling in new tasks. KlF also demonstrates strong generalizability, making it suitable for various base models and adaptable to PEFT methods like LoRA. Furthermore, it offers notable extensibility, supporting enhancements through integration with memory replay techniques. Comprehensive experiments conducted on two CL benchmarks, involving models ranging from 220M to 7B parameters, affirm the effectiveness of KlF and its variants across different settings.
@@ I Introduction
Figure 1: Conceptual illustration of KlF. By identifying task-relevant areas across both previously accumulated and current tasks, we can consolidate the task-shared and task-specific parameters to facilitate efficient knowledge transfer (KT) and mitigate catastrophic forgetting (CF).
@@
To address these limitations, we introduce Knowledge Localization and Fusion (KlF) [19], a novel CL framework designed to improve KT between tasks without relying on memory replay. Our approach is motivated by recent findings that model parameters contribute unevenly to performance [20]. For instance, the authors in [21] uncovered a core region in LLMs that is crucial for all languages, suggesting that preserving these vital regions can mitigate forgetting. Additionally, findings from [14] indicate that fine-tuning with LoRA often preserves many unnecessary parameter changes, pointing to substantial redundancies within PEFT blocks.
@@
Based on these insights, KlF facilitates KT by pinpointing and consolidating the importance distribution of model parameters across tasks. Initially, KlF utilizes a group-wise importance-aware knowledge localization technique that employs gradient trajectories to identify tiny regions within the parameter space that store crucial knowledge for the current task. By comparing the importance distribution with those of previous tasks, we can then differentiate between task-specific and task-shared regions, as illustrated in Figure 1. The subsequent knowledge fusion phase then categorically integrates weights from previous tasks with the current one, efficiently mitigating CF.
@@
In detail, KlF first reconstructs the model or PEFT block into fine-grained “skill units”. A skill unit refers to a distinct subset of model parameters that encapsulates specific functional capabilities or knowledge relevant to a particular task, such as the Query matrix within the self-attention layer. By operating at this finer granularity, we can localize and consolidate task-specific and shared knowledge within a single PEFT block, rather than adapting a separate PEFT block for each task as in previous works (addressing Limitation 1).
@@
The importance-aware knowledge localization method employs a new group-wise metric to compute importance scores, effectively quantifying the significance of each skill unit for the current task. Our knowledge fusion phase, then based on a fine-grained model averaging strategy, effectively manages different types of knowledge. This stage facilitates forward KT by using previously fine-tuned weights as the starting point for new tasks. For backward KT, we merge knowledge from both current and past tasks into localized task-shared skill units, boosting their effectiveness. To prevent CF, we ensure the integrity of skill units containing prior task-specific knowledge is maintained, safeguarding them from being altered by the learning of new tasks (addressing Limitation 2).
@@
Given the widespread adoption and success of LoRA in fine-tuning LLMs, there is a compelling opportunity to optimize the KlF framework specifically for LoRA. While the initial KlF framework has shown promising results, its validation has been limited to a specific sub-task, and its broader applicability on general CL benchmarks remains explored. Additionally, the reliance on first-order gradients for knowledge localization in KlF may not yield precise importance localization. The knowledge fusion phase also involves many hyperparameters, which could complicate the averaging process.
@@
To address these issues, we propose KlFLoRA, a variant of KlF optimized for LoRA. KlFLoRA redesigns the LoRA adapter into new skill units based on parameter dependencies, streamlining knowledge management throughout sequential task learning. It incorporates an orthogonal loss during fine-tuning to ensure distinct latent semantic features are captured within the same LoRA adapter. For knowledge localization, KlFLoRA utilizes a new second-order gradient approximate group-wise metric, enhancing the precision of importance scores. In knowledge fusion, it shifts from hard-mask to soft-mask averaging, employing an adaptive technique that flexibly integrates task-specific and shared parameters according to the importance of the skill unit. Additionally, we combine KlF with memory replay to broaden its application scope, introducing the KlF-M model. KlF framework and its variants excel in mitigating CF and showcase remarkable capabilities for KT, outperforming SOTA methods.
@@
  • We introduce the Knowledge Localization and Fusion (KlF) framework for language model continual learning. By pinpointing and amalgamating task-specific and task-shared knowledge at a granular skill unit level, KlF ensures effective knowledge transfer and significantly reduces catastrophic forgetting, addressing the shortcomings of previous methodologies.
  • We devise various group-wise knowledge localization and fine-grained knowledge fusion techniques. Notably, our KlF localization utilizes either first-order gradients or an innovative second-order gradient approximation for assessing parameter importance. For knowledge fusion, we implement strategies ranging from categorical hard-masking to adaptive soft-masking, enhancing the flexibility and precision of knowledge integration.
  • The KlF framework is distinguished by its robust generalizability and extensibility. Its skill units are designed flexibly, allowing for seamless tailoring to PEFT methods, such as LoRA, optimizing performance across different model architectures. Moreover, integrating KlF with memory replay enables further performance enhancements and broadens its applicability to diverse scenarios.
  • Extensive evaluation on two CL benchmarks demonstrates the superiority of our KlF framework and its variants in facilitating knowledge transfer and mitigating catastrophic forgetting, especially in memory-free scenarios. Furthermore, KlF consistently excels across diverse model sizes (from 220M to 7B), various model architectures (T5 and LLaMA-2), and unseen tasks.
@@ II-A Language Model Continual Learning
Conventional continual learning methods can be categorized into three main types: (i) Regularization-based methods impose explicit constraints to preserve the knowledge of previous tasks by penalizing the variation of each parameter based on its contribution to past tasks [12, 22], such as in the EWC [23] method. However, these methods typically restrict the learning of task-shared knowledge by only considering parameter importance from the perspective of past tasks. This limitation often leads to a suboptimal balance between retaining previous knowledge and excelling at new tasks, as it does not account for the relevance of parameters to both past and current tasks. (ii) Rehearsal-based methods mitigate catastrophic forgetting by either storing old training samples [24, 25] or training generative models to provide pseudo-samples of previous tasks [26, 27], enabling data replay during the learning of new tasks. (iii) Architecture-based methods aim to prevent task interference by dynamically expanding model capacity or assigning dedicated parameter subspaces to each task [28, 29]. While effective at minimizing forgetting, these methods often hinder knowledge transfer (KT) between tasks. Our KlF framework stands out by bi-directionally analyzing parameter importance across both historical and current tasks, enabling effective KT while mitigating forgetting. To further enhance performance, we extend KlF with memory replay techniques, resulting in KlF-M, improving its flexibility and applicability across various scenarios.
Language model continual learning with PEFT methods in the era of LLMs often adopt a parameter isolation strategy. This approach involves using a pipeline to learn and select PEFT blocks for each task [11, 30]. However, as the number of tasks increases, the accumulation of PEFT blocks can become inefficient, and achieving KT between blocks remains challenging. More recently, authors in [31] have explored using the Mixture of Experts (MoE) approach to connect multiple PEFT blocks, but this essentially relies on coarse-grained model averaging. This can lead to overemphasizing unimportant weights, contaminating previously acquired task-specific knowledge, and causing forgetting. To address these challenges in the context of PEFT-based CL, we extend the KlF framework and introduce KlFLoRA. By dividing fine-grained skill units based on parameter dependencies within LoRA, KlFLoRA enhances KT efficiency, enables more precise control, and effectively mitigates forgetting.
II-B Knowledge Localization
@@
Motivated by these findings, we address the challenge of catastrophic forgetting in CL. While [20] introduced the concept of “skill localization” to identify critical parameters in pre-trained language models, their approach requires additional steps to identify and retrain these parameters post-fine-tuning, which impacts efficiency. In contrast, our importance-aware knowledge localization method leverages trajectory gradients during each task’s training phase to identify parameter importance, significantly reducing processing time. Combined with our knowledge fusion strategy, which explicitly targets task-specific and shared parameters, our approach effectively mitigates forgetting and achieves knowledge transfer in CL.
@@ III-A Continual Learning Setup
And we denote $f_{k}$ as the model trained up to task $\mathcal{T}_{k}$, while $\hat{f}_{k}$ represents the model after integrating knowledge from $\hat{f}_{k-1}$ and $f_{k}$ through averaging. Our KlF framework is designed to tackle a more challenging scenario where the model cannot access any historical data during training [18]. To broaden its applicability, we have augmented KlF with memory replay. In this enhanced setup, known as KlF-M, we store a random subset of $\left|\mathcal{M}\right|$ samples from each prior task’s training set $\mathcal{D}_{i}$ in memory $\mathcal{M}_{i}$. The model is then trained jointly on the new task data $\mathcal{D}_{k}$ and the accumulated memory $\mathcal{M}_{<k}$, leveraging this memory replay to improve performance.
@@ IV-A Overview
KlF includes two key components: (i) Knowledge Localization, utilizing a group-wise importance metric to precisely identify the importance distribution of parameters across tasks, and (ii) Knowledge Fusion, which employs an innovative fine-grained model averaging strategy to integrate model weights from both current and previous tasks for effective knowledge transfer. Figure 2 presents a detailed overview of our proposed KlF framework, with the following subsections elaborating on each component and the corresponding extensions.
Figure 2: Left: Depiction of reconstructing LoRA as fine-grained skill units. Right: Overview of KlF. Step 1: We compute the importance scores of skill units for the current task $k$ using our importance-aware knowledge localization method during fine-tuning. Step 2: Based on a categorical model averaging strategy, the knowledge fusion mechanism merges the model $\hat{f}_{k-1}$, which accumulates knowledge of all previous tasks, with the current task’s model $f_{k}$. This integration is strategically guided by the importance distributions of skill units across different tasks. This iterative process continues with the addition of each new task.
@@ IV-B1 Matrix-Level Skill Unit
In this division strategy, we define skill units as individual matrices in the model, such as the query or key matrices in the self-attention layer or the A matrix in LoRA. This approach aligns with the original KlF framework and offers the advantage of being model-agnostic. It is suitable for both traditional full-parameter fine-tuning and modern parameter-efficient fine-tuning methods like LoRA, ensuring broad generalizability.
@@ IV-B2 LoRA-Tailored Skill Unit
IV-C Importance-aware Knowledge Localization
@@ IV-C1 First-Order Gradient-Based Metric
In the original KlF framework, we defined the importance function $s(\cdot)$ as the magnitude of the gradient-weight product:
@@ IV-C2 Second-Order Gradient Approximation Metric
Algorithm 1 Importance-aware Knowledge Localization
@@
where $\beta\in[0,1]$ and Norm($\cdot$) normalizes the importance scores to the [0, 1] range, ensuring consistency across models. The initial scores, $\mathcal{I}(\hat{\mathcal{U}}_{1})$, are set equal to $\mathcal{I}(\mathcal{U}_{1})$. Following this, the importance distribution for skill units up to task $\mathcal{T}_{k-1}$ is combined with that of the current task, $\mathcal{T}_{k}$, to facilitate the knowledge fusion process.
IV-D Knowledge Fusion
Following knowledge localization, the next critical phase is knowledge fusion, where knowledge within each skill unit is integrated into a cohesive framework. This process utilizes a sophisticated model averaging approach that accounts for various factors to optimize task performance.
@@ IV-D Knowledge Fusion
IV-D1 Static Weighted Fusion
In the original KlF framework, this merging strategy initiates by establishing importance thresholds $\delta$ through quantiles, selecting the top 20% of skill units based on importance scores. A skill unit $u_{i}^{k}$ is considered important (denoted as $(u_{i}^{k})^{+}$) if its score $\mathcal{I}(u_{i}^{k})$ exceeds $\delta_{k}$, and unimportant ($(u_{i}^{k})^{-}$) otherwise.
The static weighted fusion strategy tailors parameter integration for each skill unit, according to its importance across different tasks. The method is defined as follows: